Bramble Corporation had net sales of $5,060,000 and interest revenue of $86,000 during 2017. Expenses for 2017 were cost of goods sold $3,120,000; selling expenses $434,000; administrative expenses $550,000; and interest expense $23,000. Bramble’s tax rate is 40%. The corporation had 560,000 shares of common stock authorized and 56,000 shares issued and outstanding during 2017. Prepare a single-step income statement for the year ended December 31, 2017.

BRAMBLE CORPORATION Income Statement For the Year Ended December 31, 2017
Answer #1

Answer

Bramble corporation

Income statement  

For the Year Ended 31st December, 2017

Particulars $ $
Revenue:
Net sales 5,060,000
Interest revenue 86,000
Total Revenue 5,146,000
Less Expenses:
Cost of goods sold 3,120,000
Administrative expense 550,000
Interest expense 23,000
Selling expense 434,000
Tax expense 407,600
Total expenses 4,534,600
Net income 611,400
Earnings per share 10.92

Calculations

Earnings before tax

= 5,146,000 - (3,120,000 + 550,000 + 23,000 + 434,000)

= 5,146,000 - 4,127,000 = $1,019,000

Tax rate is 40%

Income tax expenses = 1,019,000 × 40% = $407,600

Income tax expense is $407,600.

Earnings per share = Income available to common share holders / weighted average number of common shares outstanding

56,000 shares issued

Earnings per share = 611,400 / 56,000 = $10.92

$10.92 is earnings per share.

Single step income statement and its related calculations are presented above.

If YOU are like the answer please upvote.